OpenStudy (anonymous):

What is the nth term of the geometric sequence that has a common ratio of 6 and 24 as its third term? an=24(6)n−1 an=23(6)n−1 an=24(6)n an=32(6)n−1

OpenStudy (cwrw238):

third term = a1 * r^2 = a1* 6^2 = 36 * a1 36 a1 = 24 a1 = 2/3 nth term = (2/3) * 6^(n-1)